### 22 Jan 2024 v1

# 3cm3cm2cm2cm Checking the Be anomaly with a two-arm electron positron pair spectrometer

Abstract
We have repeated the experiment performed recently by Krasznahorkay et al., (Phys. Rev. Lett. 116, 042501 (2016)), which may indicate a new particle called X17 in the literature. In order to get a reliable, and independent result, we used a different type of electron-positron pair spectrometer which have a more simple acceptance/efficiency as a function of the correlation angle, but the other conditions of the experiment were very similar to the published ones. We could confirm the presence of the anomaly measured at the E=18.15 MeV resonance, and also confirm their absence at the E=17.6 MeV resonance, and at E= 800 keV off resonance energies.
1 Introduction
An experiment was conducted at the ATOMKI Laboratory (Debrecen, Hungary) [1] in 2016, studying the Li(p,ee)Be nuclear reaction. The target nucleus was excited through proton capture, with the experiment set up to detect ee pairs produced in Internal Pair Creation (IPC) during the transition from the excited to the ground state of Be. The experimental setup was using a set of multiwire proportional counters placed in front of E and E detectors, to determine the ee opening angle, . The very thin E detectors were made of plastic scintillators and chosen to provide excellent gamma suppression. While the much thicker E detectors were used to measure the total energy of the electron and positron. A detailed description of the experimental setup can be found in [2]. The ATOMKI collaboration observed a deviation in the distribution with respect to the expected Rose theory distribution [3], at around 140 degrees.
After 2016 the ATOMKI collaboration repeated the measurements, while improving the experimental setup in many different ways [4, 5]. The anomaly kept appearing in the follow-up experiments, with no nuclear physics model being able to explain it. This led to the explanation of a new particle, beyond the Standard Model of particle physics, created and decaying to ee pairs, being detected in these experiments. The hypothetical particle is now commonly referred to as X17, because of the invariant mass calculated from the e+e- anomalies.
In addition to the Li(p,ee)Be nuclear reaction, since 2019 the same collaboration has also studied the H(p,ee)He and B(p,ee)C reactions. Different proton energies were used, leading to anomalies appearing in the distributions at different angles. All well consistent with the assumption of an 17 MeV particle being created with different kinetic energies, leading to different opening angles between the electron positron pairs.
The ATOMKI anomaly being a genuine physics effect is supported by a number of arguments.
The anomaly has been observed in Be with experimental setups using different geometries, with 5 and 6 arm spectrometers.
The anomaly has been observed using fundamentally different position sensitive detectors: multi-wire chambers and silicon strip detectors.
The anomaly has been observed in three different nuclei by now (Be, He and C), showing up at ee opening angles consistent with a single particle.
The anomaly has been observed at different proton beam energies at varying ee opening angles, also consistent with a single particle.
All observed anomalies have a very high statistical significance.
Despite the consistency of the numerous observations at ATOMKI, more experimental data is needed to understand the nature of this anomaly. Many experiments around the world started looking for such a particle in different channels, or are planning to do so. Many of these experiments already put constraints on the coupling of such a hypothetical particle with ordinary matter. Others are still in an R&D phase, soon to contribute to a deeper understanding of this phenomenon as concluded by the community report of the Frascati conference [6]. Such report gives also a nice overview of the possible theoretical interpretation of the observed anomalies in Be, He and C [6].
The aim of this article is to look for the Be anomaly at the VNU University of Science (HUS) with a two-arm electron-positron spectrometer specifically designed, and built for this purpose.
2 Experimental methods
In order to make our experimental results easier to reproduce, in the first part of this paper we give a more precise description of our pair spectrometer with the setup of the detectors and the electronics connected to them. This will be followed by a brief description of the data acquisition system and the results of testing and calibrating the detectors.
2.1 The ee spectrometer
In the present experiment two detector telescopes including Double Silicon Strip Detectors (DSSD) and plastic scintillators were used, placed at an angle of 140 with respect to each other. The diameter of the carbon fiber tube of the target chamber has been reduced from 70 mm (used in ATOMKI) to 48 mm to allow a closer placement of the telescopes to the target. This way we could cover a similar solid angle at 140 like the one used in the ATOMKI experiment Fig 1.
In this setup the efficiency function has only one maximum as a function of the opening angle. This angular dependence can be simulated and calibrated more robustly than fore more complicated configurations. Another advantage of this setup that we used is that its sensitivity to background from cosmic radiation is significantly less.
The two detector telescopes were placed at azimuthal angles -20 and -160 with respect to horizontal. This meant that cosmic rays, predominantly arriving vertically, would have a very low chance of hitting both telescopes at the same time.
For the detection of 1-20 MeV e and e particles, which are considered high-energy in nuclear physics, GEANT4[7] calculations were performed for different detector materials. This included plastic scintillators, Ge semiconductor detectors and LaBr scintillators. Much better energy resolution could be achieved with the latter two types of detectors than with of a plastic scintillator. However, our simulations also showed that in the case of high-density and high Z detectors, the efficiency of the total energy detection in the energy response function is greatly reduced compared to the integral of the response function. In materials with higher atomic numbers, the e and e particles slow down faster, and thus the probability of generating bremsstrahlung radiation is higher. However, there is a high probability that these radiations will escape the detector. This is why the probability of detecting e and e particles at full energy is reduced in these high Z materials. Therefore, we used special plastic scintillators for this task. The dimensions of the EJ200 plastic scintillators were chosen (82 × 82 × 80 mm each) in such a way that these high-energy particles would be completely stopped in them.
To collect the light generated by the ee particles slowing down in the detector from all corners with the same efficiency, the SCIONIX company provided the detector with specially shaped light guides. The surfaces of both the detectors and the light guides were diamond polished. The collected light was converted to an electronic signal by Hamamatsu PMT assemblies were used.
The time resolution of the detector turned out to be adequate despite the large size of the detector and multiple light reflections inside. It was measured with a Co source between 2 detectors, and found to be less than 1 ns Fig 2.
In the absence of a high-energy electron source, the energy resolution of the detector could only be determined with the help of ee pairs from internal pair creation. We will come back to this in section 2.4 in connection with the energy calibration of the detectors.
Double-sided Silicon Strip Detectors (DSSD) were used to measure the energy loss of the particles and their directions. The DSSDs were purchased from Micron Semiconductor. They consist of 16 sensitive strips on the junction side and 16 orthogonal strips on the ohmic side. Their element pitch is 3.01 mm for a total coverage of 49.5 x 49.5 mm. They are mounted on a PCB board, with 34 pins on one edge for their readout. These are connected via 34-conductor flat cables to the pre-amp boxes.
Mesytec MUX32 type 32-channel preamplifiers, linear amplifiers, timing filter amplifiers, timing discriminators and multiplexers were used. MUX32 is a very fast 2x16 channel multiplexed preamplifier, shaper and discriminator combination with very good energy and timing resolutions. Up to two simultaneously responding channels are identified and two amplitudes plus the two corresponding amplitude coded addresses (position signals) are sent to the outputs. These modules are especially well suited for DSSD detectors.
By properly shielding and grounding the detectors, using 10 m thick Al foils mounted on PCB boards both at the front and the back side of the detectors and shielding the flat cables and connectors, we managed to reduce their electronic noise and could lower the levels of the discriminators below 50 keV. For simplicity, the detectors are operated in air without dedicated cooling. The first test and calibration of the detectors was performed with a mixed -source containing Pu, Cm, and Am. The energy loss in the Al foil and in the air (5 mm) were taken into account.
2.2 Data acquisition system
The data acquisition system uses VME ADC, TDC, and QDC units, which were read out with the help of a commodity desktop PC. The software used for configuring the VME devices and recording the data collected can be found in [8].
CFD discriminators and TDC units were used to determine the arrival time of the signals coming from the plastic scintillators, and QDCs were used to digitize their energy signals. The block diagram of the electronics connected to the detectors is shown in Fig 3. The signals received from the MUX32 multiplexer were digitized with the help of ADC and TDC units.
2.3 Calibration of the DSSD detectors
Energy of the DSSD were calibrated using information from the ee particles, which were produced by the Li(,)Be nuclear reaction. The energy of the bombarding protons was set to the E=441 keV resonance. When creating the spectrum, we required real coincidence between the DSSD detector and the plastic scintillator located behind it, which measured the total energy of the ee particles. We also required this energy to be in the 6-20 MeV range. The experimental configuration was used in GEANT4 simulation code[7]. Fig 4 shows the energy loss distribution of electrons and positrons passing through one of the DSSD detectors.
The histograms shows reasonably good agreement between the experimental and simulated distributions, however we obtained some differences around 100 keV and 350 keV. The former can be caused by some electronic noise, while the latter is caused by the events when both members of the ee pair created during internal pairing passed through the same DSSD detector. The detection efficiency with a CFD threshold of 50 keV was greater than 97% for electrons and positrons.
The histograms in Fig 5 show the (non calibrated) distribution of the x and y coordinates of the impact points of the e/e particles that hit the DSSD detector. The peaks of the fence spectrum correspond to the coordinates of particles passing through the individual silicon strips. As shown, the assignment of the recorded data to the x, y strips is very clear. The width of DSSD was used to convert the recorded data to position in cm scale with origin at the center (two dimension plot in Fig 5). Of the 70,000 events recorded for these distributions, only a few had to be excluded.
2.4 Calibration of the scintillation detectors
The energy distribution of the electrons and positrons in internal pair creation individually is contiguous. However if the electron and positron lose their energy in the same detector, the energy distribution of such events will show peaks at the transition energies - . The scintillator detectors were calibrated using such events. If the detectors are close enough to the target there is a high probability of the above internal energy summing.
The energy spectrum of the scintillators, measured by the scintillators for events selected in this way (by gating on “multiplicity=2” events in the DSSD detector), can be found in Fig. 6.
3 Experimental results
The experiments were performed in Hanoi (Vietnam) at the 1.7 MV Tandem accelerator of HUS, with different proton beam energies between 0.4.
LiF targets with thicknesses of 30 evaporated onto 10 m thick Al foil, as well as LiO targets with thicknesses of 0.3 mg/cm were used on 1 m thick Ni foils in order to maximize the yield of the pairs.
radiations were detected by a 3”x3” NaI(Tl) detector monitoring also any potential target losses. The detector was placed at a distance of 25 cm from the target at an angle of 90 degrees to the beam direction.
A single energy spectrum measured by the scintillators and gated by “multiplicity=2” events in the DSSD detector, which means that both the electron and positron coming from the internal pair creation are detected in the same telescope, is shown in Fig. 6 for telescope 1. The figure clearly shows the transitions from the decay of the 17.6 MeV resonance state to the ground and first excited states in Be. The cosmic ray background was already subtracted from from that.
In order to check the effective thickness of the targets, we measured the excitation function of Be via Li(p,)Be reaction by scanning the proton beam energies from 441 keV to 1300 keV. The events with multiplicity-2 in DSSD and measured energy in plastic larger than 10 MeV were counted, and their rate was plotted in Fig 7.
There are two resonance peaks were observed at the proton beam energies of round 441 and 1030 keV[9]. The width of the resonance shows the effect of the target thickness.
The efficiency (acceptance) as a function of the correlation angle in comparison to isotropic emission was determined from the same data set by using uncorrelated ee pairs formed of separate, single events[2]. To do this, uncorrelated ee pairs have been recorded during the experiment. The analysis would select event pairs with uncorrelated electrons/positors hitting different telescopes in the events. The opening angle distribution of electron/positron pairs from such events is shown by the black line in Fig 8 (left). The green line is a simulation curve produced using Geant4. They show quite a good agreement between the estimated experimental and simulation efficiencies.
Coincidence events, with both arms of the spectrometer detecting e/e particles, were also recorded. The opening angle distribution of ee pairs from such events is shown in Fig 8 (right).
In the first experiment, we used a proton beam energy of 411 keV to bombard the Li target. Through this, the Be nucleus would be created in the 17.6 MeV excited state. Figure 9 shows the angular correlations of ee pairs originating from the transition of this 17.6 MeV excited state of Be to its ground state.
The Monte Carlo detector simulations of the experiment were done using Geant4 and are shown as histograms in Figure 9 for M1 (cyan) and E1 (green) multipolarity transitions. The simulation included the geometries of the target chamber, target backing and detector arm assemblies. The interaction of generated electrons, positrons and gamma rays were then simulated with the experimental setup. Internal Pair Creation (IPC) events, generated from both of the possible E1 and M1 transitions, were simulated this way. The combination of the E1+M1 distributions show a good agreement with the experimental data.
In the second experiment, we changed the proton beam energy to 800 keV at the off-resonance energies. With the same method to build the total simulation curve and show the experimental data in the Fig 10 we can see the simulated curve go through the middle of data points. There is no systematic deviation from the experimental points from the IPC simulation curve.
Finally, we changed the proton beam energy to the second resonance peak shown in Figure 7 (E = 1.04 MeV). The combined IPC simulation curve and experimental data for this transition is shown in Fig 11. It clearly shows a deviation in the ee opening angle distribution between the data and the simulation around . This deviation is in agreement with the result published by the ATOMKI collaboration in [1]. By assuming that the deviation is coming from the creation and immediate decay of an intermediate particle to an ee pair, we can calculate a mass of (stat.) MeV for this particle. With a deviation of 4-5 sigma from the Standard Model hypothesis.
The difference in the angle of the anomaly obtained in this study of , and the published one of [1], could be explained by a slight difference in the proton beam spot position on the target. A shift of just 0.5 mm could explain this difference. The systematic uncertainty on the calculated particle mass from the beam spot’s position was estimated using a series of simulations, using different beam spot positions. This resulted in a MeV uncertainty.
4 Summary
We successfully built a two-arm ee spectrometer in Hanoi. The spectrometer was tested and calibrated using the 17.6 MeV M1 transition excited in the Li(p,ee)Be reaction. We have got a nice agreement between the experimentally determined acceptance of the spectrometer with the one coming from our simulation. The angular correlation of the ee pairs measured for the 17.6 MeV transition agrees well with the simulated one for the M1 transition. No anomaly was observed for this, or the = 800 keV transition either. However, a significant anomaly () was observed for the 18.15 MeV transition at around , in agreement with the ATOMKI results published in 2016[1]. The mass of the hypothetical particle from this result was obtained to be (statistical) (systematic) MeV. In the future, we are planning to upgrade the spectrometer to get a wider angular acceptance.